# Managing Classical Processing Requirements for Quantum Error Correction

### v1

@@ Abstract.
Quantum Error Correction requires decoders to process syndromes generated by the error-correction circuits. These decoders must process syndromes faster than they are being generated to prevent a backlog of undecoded syndromes that can exponentially increase the memory and time required to execute the program. This has resulted in the development of fast hardware decoders that accelerate decoding. Applications utilizing error-corrected quantum computers will require hundreds to thousands of logical qubits and provisioning a hardware decoder for every logical qubit can be very costly. In this work, we present a framework to reduce the number of hardware decoders and navigate the compute-memory trade-offs without sacrificing the performance or reliability of program execution. Through workload-centric characterizations, we propose efficient decoder scheduling policies which can reduce the number of hardware decoders required to run a program by up to 10$\times$ while consuming less than 100 MB of memory.
@@ 1. Introduction
As quantum computing enters a phase of rapid scaling to enable Fault-Tolerant Quantum Computing (FTQC), the classical processing resources required to support Quantum Error Correcting (QEC) codes must be scaled proportionally. QEC codes generate a stream of syndromes repeatedly by measuring parity qubits every cycle, and a decoder algorithm running on the classical control computer processes the stream of syndrome bits to detect errors and correct errors. Recent demonstrations have shown how the Surface Code (Fowler2012, ) can be deployed experimentally to suppress logical error rates (Google2023, ), how neutral atoms can be used to realize up to 48 logical qubits (Bluvstein2023, ), and how four logical qubits could be created with thirty physical qubits to achieve an 800$\times$ reduction in the error rate (quantinuum2024, ). These demonstrations are precursors to complex systems with more logical qubits requiring significant classical processing resources to enable fault-tolerant architectures.
@@
Applications that can benefit from FTQC will require hundreds to thousands of logical qubits to function (Blunt2024, ). Depending on the error-correcting code used, replicating decoders for every logical qubit in the system can become very expensive and intractable in terms of cost and complexity. To reduce this cost, fast, hardware-efficient decoders have been proposed which sacrifice some accuracy for speed and scalability by making approximations in the decoding process (Ravi2023, ; Riverlane2023, ; Vittal2023, ; Alavisamani2024, ). However, catering to hundreds to thousands of logical qubits with these specialized decoders will still result in complex and costly systems – in this work, we aim to show how the total number of decoders can be reduced without affecting the performance or reliability of the quantum computer, thus allowing for more scalable classical processing for QEC.
Figure 1. (a) The compute and memory trade-offs for the classical processing required to implement QEC – fewer decoders result in an exponential increase in the memory required to store undecoded syndromes, and we aim to avoid this exponential growth in memory while reducing the number of decoders; (b) Virtual decoders – a scheduling policy will allocate decoders to the logical qubits.
@@
In this paper, we present VQD: Virtual Quantum Decoding, a framework that aims to provide the illusion that there are decoders for every logical qubit while using significantly fewer hardware decoders to enable scalable and efficient classical processing necessary for fault-tolerant quantum computers. As shown in Fig. 1(a), the objective of VaDER is to reduce classical computing and memory resources needed to execute quantum programs on a fault-tolerant quantum computer without causing a performance slowdown or increase in the logical error rate. This is challenging because when we reduce the number of physical decoders, the memory required to store undecoded syndromes can grow exponentially if the syndrome generation and syndrome processing rates are not matched. More importantly, this exponential increase in memory due to syndrome backlog will lead to an exponential increase in the time it takes to process all the undecoded syndromes, thereby significantly increasing the execution time of the program (Terhal2015, ). Our experimental evaluations also show that if a logical qubit is not decoded for extended periods (which will occur if there are fewer decoders than qubits), then it can cause the decoder latency to increase due to an increase in undecoded errors. This increase in the decoder latency can affect the application of non-Clifford states – if decoding is delayed for a logical qubit before applying a non-Clifford state, the application of the non-Clifford state could be delayed since the decoder might take more time than usual to decode all prior rounds.
@@
Given the challenges in sharing decoder hardware and to understand how it can be enabled, we characterized representative FTQC workloads to understand the decoding requirements from a performance and reliability perspective. Our characterization using a lattice surgery compiler (watkins2023high, ) revealed that there is a limited amount of operational parallelism due to long sequences of $T$ and $H$ gates resulting from Clifford + $T$ decomposition, which is necessary for universality. This is crucial, as non-Clifford gates are the reason real-time decoding is necessary for FTQC. Non-Clifford operations are the only operations where the decoding is in the critical path, and fortunately, they occur in a highly serialized manner. Therefore, a physical decoder per logical qubit is unnecessary and will lead to severe underutilization.
@@
Armed with this insight, we propose a system architecture with significantly fewer physical decoders than the number of logical qubits. Furthermore, we design efficient decoder scheduling policies for such systems. Such a system can be visualized in Fig. 1(b). We propose a scheduling policy that minimizes the Longest Undecoded Sequence, termed as the MLS policy. We compare it with the Round Robin (RR) and Most Frequently Decoded (MFD) policies – our evaluations show that the MLS policy can reduce the number of hardware decoders by up to 10$\times$ while ensuring that no logical qubit remains undecoded for a significantly long period.
We also propose a noise-adaptive scheduling policy that can prioritize decoding of logical qubits that incur a sharp increase in the physical error rate due to phenomena such as cosmic rays (McEwen2021, ) and leakage due to heating (Miao2023, ). This involves a simple detector that can schedule decoding for a logical qubit in case the syndromes for that logical qubit show a sudden increase in bit-flips. Next, we show how some decoding tasks can be offloaded to software to further improve the efficacy of decoder scheduling policies.
Balancing compute and memory is a classic architectural problem, and we use VQD to explore these trade-offs. Prior research on decoders and classical processing required for fault-tolerant quantum computers have focused on reducing the hardware cost of implementing decoders by making approximations in the decoding algorithm, sometimes at the cost of accuracy (Vittal2023, ; Riverlane2023, ; Alavisamani2024, ). With VaDER, we show that even if individual decoders have a high hardware cost, the overall cost can be reduced significantly by virtualizing decoders.
@@ 2.1. Quantum Error Correction
Quantum Error Correction (QEC) improves the reliability of a system by utilizing many physical qubits to encode a single logical qubit (Shor1995, ; Knill1997, ). Most QEC codes can be categorized as stabilizer codes (Terhal2015, ) – some promising stabilizer codes include quantum Low Distance Parity Check (qLDPC) codes (Bravyi2024, ) and Surface Codes (Fowler2012, ). Owing to their relatively relaxed connectivity requirements that can be realized with hardware available today, we focus specifically on the Surface Code. Note that this work can be extended to other QEC codes apart from the Surface Code.
@@
A single rotated Surface Code patch of distance $d=3$ is shown in Fig. 2(a) (Horsman2012, ). Black circles denote data qubits and each data qubit is connected to $X$ and $Z$ parity qubits. The Surface Code works by repeatedly measuring syndromes, which correspond to the measurements performed on all $X$ and $Z$ measure qubits in a patch after performing the sequence of gates shown in Fig. 2(b). By repeatedly measuring these syndromes, both bit-flip and phase-flip errors occurring within a patch can be detected by the decoder.
@@ 2.2. Decoding Errors
Fig. 2(c) shows the general procedure of how any generic QEC code works – syndromes are constantly being generated, which are then fed to a decoder. Syndromes contain information about which qubits have flipped in every round of syndrome measurements, and these flips allow the decoder to determine what errors on the data qubits caused those flips. Since errors can always be expressed in the form of Pauli gate, they can be corrected in software without executing any physical operations on the logical qubits. This is achieved by updating the Pauli frames for all data qubits that make up that logical qubit, which adjusts the interpretation of future measurements by accounting for the error that was detected (Terhal2015, ). For the Surface Code, the decoding problem is commonly formulated as a Minimum Weight Perfect Matching (MWPM) problem, which leverages a graph representation of the syndrome measurements (Fowler2015, ; Wu2022, ).
@@
2.3. Non-Clifford Gates
On a Surface Code error-corrected quantum computer, all Clifford gates can be performed reliably either in software or via logical operations performed via Braiding (Fowler2012, ) or Lattice Surgery (Horsman2012, ). However, non-Clifford gates such as the $T$ gate cannot be applied in a fault-tolerant manner directly. This is because logical qubits initialized with the $T$ gate will have an error probability equal to the underlying physical error rate of the system, $p$, thus making them impure (Litinski2019magic, ). However, multiple impure states can be used to distill fewer, purer logical qubits with a $T$ state (known as a magic state $|m\rangle$) – this process is known as magic state distillation (Bravyi2012, ; Gupta2024, ).
Fig. 3 shows how a $T$ gate can be applied to a logical qubit $P$ in a fault-tolerant manner by using Lattice Surgery (LS) (Litinski2019, ). The magic state $|m\rangle$ is a purified $T$-state. Since a non-Clifford gate is being applied, all prior errors that affected $P$ must be known before $|m\rangle$ is applied to prevent errors from spreading (Terhal2015, ). Lattice Surgery can be used to perform a $Z\otimes Z$ operation on $P$ and $|m\rangle$ to apply the magic state (Litinski2019, ; Litinski2019magic, ). Once the $Z\otimes Z$ operation is performed, the decoding result of $P$ prior to Lattice Surgery is combined with the decoding result of Lattice Surgery multi-body measurement and the measurement of the patch containing the magic state $|m\rangle$ to determine an appropriate Clifford correction [^1] The auto-corrected $\pi/8$ gate in (Litinski2019, ; Litinski2019magic, ) uses an additional ancillary qubit that has not been shown.. This correction needs to be known before the next logical operation involving a non-Clifford gate.
[…]
[some figure]
Figure 4. (a) Exponential increase in the memory required to store undecoded syndromes for different number of available decoders for the wstate-60 benchmark – 200 decoders correspond to assigning a decoder to every logical qubit; (b) Number of $T$-gates required for different workloads; (c) Exponential increase in the decoder latency per round (in nanoseconds) as the number of rounds is increased from $d$ to $20d$ – the increase is higher for larger code distances ($p=10^{-4}$); (d) Slow increase in the logical error rate as the number of rounds of error correction increase with $p=10^{-3}$, $p=10^{-4}$; (e) Histogram of the number of concurrent critical decodes for different workloads with the EDPC layout – there is limited parallelism as far as critical decodes are concerned (f) Number of FPGAs required for decoders to implement different workloads with (i) a Single $T$-factory and (ii) Optimal number of $T$-factories.
@@ 2.4. Critical Decodes
For a logical qubit that is only executing Clifford gates, errors can be decoded at any point of time (even after the experiment has ended). This is because all Clifford corrections can be commuted to the end of the circuit, essentially allowing the syndromes to be post-processed rather than decoded in real-time (Terhal2015, ). However, universal fault-tolerant quantum computer require non-Clifford gates such as the $T$-gate – this makes real-time decoding a necessity since syndromes for a patch must be decoded before the application of a non-Clifford gate. We call decodes that must happen before the application of a non-Clifford gate critical decodes since all syndromes generated up to that point for that logical qubit must be decoded before computation can proceed.
@@ 3. Classical Processing Requirements
Having explained the relevant details about QEC and the role of decoders and critical decodes, we now cover the classical processing requirements for FTQC.
@@ 3.1. Syndrome Generation and Processing Rates
As discussed in Section 2, applying non-Clifford gates requires decoders to be up-to date with the latest syndrome for a logical qubit before computation can proceed. As shown by Terhal (Terhal2015, , p. 20), if the rate at which syndromes are generated $r_{gen}$ is faster than the rate at which they are processed $r_{proc}$, the memory required for storing syndromes that are yet to be decoded increases exponentially (referred to as the backlog problem). This exponential increase in the memory required also leads to an exponential increase in the runtime of the workload. Fig. 4(a) shows the exponential increase in the memory required to store undecoded syndromes for the wstate-60 benchmark with the number of rounds for various numbers of available decoders. For larger and longer running workloads, the memory requirements will be much higher. By reducing the number of available decoders, $r_{proc}$ is effectively reduced, leading to the exponential growth in memory. Decoding is necessary at runtime when consuming $T$-states, as discussed in Section 2, and Fig. 4(b) shows that it will be a frequent operation for most workloads.
$r_{proc}$ must be consistently greater than $r_{gen}$ to prevent an exponential increase in memory requirements.
3.2. Why is Real-Time Decoding Needed?
[…]
Leaving syndromes undecoded for even tens of rounds can result in an exponential slowdown in the decoder processing rate.

### v2

@@ Abstract.
Quantum Error Correction requires decoders to process syndromes generated by the error-correction circuits. These decoders must process syndromes faster than they are being generated to prevent a backlog of undecoded syndromes. This backlog can exponentially increase the time required to execute the program, which has resulted in the development of fast hardware decoders that accelerate decoding. Applications utilizing error-corrected quantum computers will require hundreds to thousands of logical qubits and provisioning a hardware decoder for every logical qubit can be very costly. In this work, we present a framework to reduce the number of hardware decoders and navigate the compute-performace trade-offs without sacrificing the performance or reliability of program execution. Through workload-centric characterizations performed by our framework, we propose efficient decoder scheduling policies that can reduce the number of hardware decoders required to run a program by up to 10$\times$. With the proposed framework, scalability can be achieved via decoder virtualization, and individual decoders can be built to maximize accuracy and performance.
Virtualization, Quantum Error Correction, Decoders
conference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japancopyright: none
@@ 1. Introduction
As quantum computing enters a phase of rapid scaling to enable Fault-Tolerant Quantum Computing (FTQC), the classical processing resources required to support Quantum Error Correcting (QEC) codes must be scaled proportionally. QEC codes generate a stream of syndromes repeatedly by measuring parity qubits every cycle, and a decoder algorithm running on the classical control computer processes the stream of syndrome bits to detect and correct errors. Recent demonstrations have shown the Surface Code (Fowler et al., 2012) can be deployed experimentally to suppress logical error rates (Acharya et al., 2024, 2023), creating 48 logical qubits using neutral atoms (Bluvstein et al., 2023), an 800$\times$ reduction in the error rate using four logical qubits (da Silva et al., 2024). These demonstrations are precursors to complex systems with more logical qubits that will require significant classical resources to enable FTQC systems.
@@
Applications that can benefit from FTQC will require hundreds to thousands of logical qubits to function (Blunt et al., 2024). Depending on the error-correcting code used, replicating decoders for every logical qubit in the system can become very expensive and intractable in terms of cost and complexity. This is especially because decoders add to the classical hardware already needed for qubit control and readout. To reduce this cost, fast, hardware-efficient decoders have been proposed which sacrifice some accuracy for speed and scalability (Ravi et al., 2023; Barber et al., 2023; Vittal et al., 2023; Alavisamani et al., 2024). However, catering to hundreds to thousands of logical qubits with these specialized decoders will still result in complex and costly systems – in this work, we aim to show how the total number of decoders can be reduced without affecting the performance or reliability of the quantum computer, thus allowing for more scalable classical processing for QEC.
Figure 1. (a) Fewer decoders result in a slowdown in the computation due to large backlogs of undecoded syndromes. VQD avoids this slowdown in computation while reducing the number of decoders; (b) Multi-objective optimization problem for designing decoders without any virtualization – current decoders are designed to minimize latency and maximize scalability, accuracy; (c) Virtualization allows the optimization problem to be reduced to just speed and accuracy – decoder virtualization can improve scalability.
@@
In this paper, we present VQD: Virtual Quantum and Decoding, a framework that aims to provide the illusion that there are decoders for every logical qubit while using significantly fewer hardware decoders to enable scalable and efficient classical processing necessary for fault-tolerant quantum computers. Reducing the number of decoders is not straightforward – if there are fewer decoders than logical qubits, some logical qubits will remain undecoded for some periods of time. When a logical qubit remains undecoded for some rounds of syndrome generation until a non-Clifford gate is executed on it, the overall decoding task is more than the original set of undecoded syndromes since syndromes keep getting generated as the decoder processes the original set of undecoded syndromes. This introduces a slowdown in the computation, as shown in Figure 1(a). The goal of this work is to avoid the slowdown in computation while reducing the compute resources required for decoding. Our experimental evaluations also show that if a logical qubit is not decoded for extended periods (which will occur if there are fewer decoders than qubits), then it can cause the decoder latency to increase due to an increase in undecoded errors. This increase in the decoder latency can affect the application of non-Clifford states – if decoding is delayed for a logical qubit before applying a non-Clifford state, the application of the non-Clifford state could be delayed since the decoder might take more time than usual to decode all prior rounds.
@@
Given the challenges in sharing decoder hardware and to understand how it can be enabled, we characterized representative FTQC workloads to understand the decoding requirements from a performance and reliability perspective. Our characterization using a lattice surgery compiler (Watkins et al., 2024) revealed that there is a limited amount of operational parallelism due to long sequences of $T$ and $H$ gates resulting from Clifford + $T$ decomposition, which is necessary for universality. This is crucial, as non-Clifford gates are the reason why real-time decoding is necessary for FTQC (Terhal, 2015). Non-Clifford operations are the only operations where the decoding is in the critical path, and fortunately, they occur in a highly serialized manner. Therefore, a physical decoder per logical qubit is unnecessary and will lead to severe underutilization.
@@
Armed with this insight, we propose a system architecture with significantly fewer physical decoders than the number of logical qubits. Furthermore, we design efficient decoder scheduling policies for such systems. We propose a scheduling policy that minimizes the Longest Undecoded Sequence, termed as the MLS policy. We compare it with the Round Robin (RR) and Most Frequently Decoded (MFD) policies – our evaluations show that the MLS policy can reduce the number of hardware decoders by up to 10$\times$ for algorithmic logical qubits and 4$\times$ when magic state distillation factories are also considered, while ensuring that no logical qubit remains undecoded for a significantly long period. We show that offloading decoding tasks to software also reduce the longest period for which a logical qubit remains undecoded by up to 40%. We also propose a dynamic scheduling mechanism that can prioritize decoding for logical qubits affected by error bursts due to phenomena such as cosmic rays (McEwen et al., 2021) and leakage due to heating (Miao et al., 2023).
We also analyze the effect of decoding latencies on the efficacy of virtualization. Our studies show that slow decoders are not amenable to virtualization, and QEC codes such as qLDPC codes will require significantly faster decoders than that are available today for virtualization to be possible. Through VQD, we show that the design space for decoders can be reduced from optimizing for latency, accuracy, and scalability (Figure 1(b)) to just latency and accuracy (Figure 1(c)). With VQD, we show that even if individual decoders have a high hardware cost, the overall cost can be reduced significantly by virtualizing decoders.
@@ 2.1. Quantum Error Correction
Quantum Error Correction (QEC) improves the reliability of a system by utilizing many physical qubits to encode a single logical qubit (Shor, 1995; Knill and Laflamme, 1997). Most QEC codes can be categorized as stabilizer codes (Terhal, 2015) – some promising stabilizer codes include quantum Low Distance Parity Check (qLDPC) codes (Bravyi et al., 2024) and Surface Codes (Fowler et al., 2012). Owing to their relatively relaxed connectivity requirements that can be realized with hardware available today, we focus specifically on the Surface Code.
@@
A single rotated Surface Code patch of distance $d=3$ is shown in Figure 2(a) (Horsman et al., 2012). Black circles denote data qubits and each data qubit is connected to $X$ and $Z$ parity qubits. The Surface Code works by repeatedly measuring syndromes, which correspond to the measurements performed on all $X$ and $Z$ measure qubits in a patch after performing the sequence of gates shown in Figure 2(b). By repeatedly measuring these syndromes, both bit-flip and phase-flip errors occurring within a patch can be detected by the decoder.
@@ 2.2. Decoding Errors
Figure 2(c) shows the general procedure of how any generic QEC code works – syndromes are constantly being generated, which are then fed to a decoder. Syndromes contain information about which qubits have flipped in every round of syndrome measurements, and these flips allow the decoder to determine what errors on the data qubits caused those flips. Since errors can always be expressed in the form of Pauli gate, they can be corrected in software without executing any physical operations on the logical qubits. This is achieved by updating the Pauli frames for all data qubits that make up that logical qubit, which adjusts the interpretation of future measurements by accounting for the error that was detected (Terhal, 2015). For the Surface Code, the decoding problem is commonly formulated as a Minimum Weight Perfect Matching (MWPM) problem (Fowler, 2015; Wu et al., 2022).
2.3. Non-Clifford Gates
On a Surface Code error-corrected quantum computer, all Clifford gates can be performed reliably either in software or via logical operations performed via Braiding (Fowler et al., 2012) or Lattice Surgery (Horsman et al., 2012). However, non-Clifford gates such as the $T$ gate cannot be applied in a fault-tolerant manner directly. This is because logical qubits initialized with the $T$ gate will have an error probability equal to the underlying physical error rate of the system, $p$, thus making them impure (Litinski, 2019b). However, multiple impure states can be used to distill fewer, purer logical qubits with a $T$ state (known as a magic state $|m\rangle$) – this process is known as magic state distillation (Bravyi and Haah, 2012; Gupta et al., 2024).
@@ 2.3. Non-Clifford Gates
Figure 3 shows how a $T$ gate can be applied to a logical qubit $P$ in a fault-tolerant manner by using Lattice Surgery (LS) (Litinski, 2019a). The magic state $|m\rangle$ is a purified $T$-state. Since a non-Clifford gate is being applied, all prior errors that affected $P$ must be known before $|m\rangle$ is applied to prevent errors from spreading (Terhal, 2015). Lattice Surgery can be used to perform a $Z\otimes Z$ operation on $P$ and $|m\rangle$ to apply the magic state (Litinski, 2019a, b). Once the $Z\otimes Z$ operation is performed, the decoding result of $P$ prior to Lattice Surgery is combined with the decoding result of Lattice Surgery multi-body measurement and the measurement of the patch containing the magic state $|m\rangle$ to determine an appropriate Clifford correction [^1] The auto-corrected $\pi/8$ gate in (Litinski, 2019a, b) uses an additional ancillary qubit that has not been shown.. This correction needs to be known before the next logical operation involving a non-Clifford gate. We assume the use of auto-corrected Pauli-Product Rotations (PPR) (Litinski, 2019b) for all non-Clifford gates.
Figure 4. (a) Exponential increase in the decoder latency per round (in nanoseconds) as the number of rounds is increased from $d$ to $20d$ – the increase is higher for larger code distances ($p=10^{-4}$); (b) Slow increase in the logical error rate as the number of rounds of error correction increase with $p=10^{-3}$, $p=10^{-4}$; (c) Histogram of the number of concurrent critical decodes for different workloads with the EDPC layout – there is limited parallelism as far as critical decodes are concerned; (d) Estimated total decoders required for different workloads with the EDPC layout and 24 distillation factories for every workload.
@@ 2.4. Critical Decodes
For a logical qubit that is only executing Clifford gates, errors can be decoded at any point in time (even after the experiment has ended). This is because all Clifford corrections can be commuted to the end of the circuit, essentially allowing the syndromes to be post-processed rather than decoded in real-time (Terhal, 2015). However, universal fault-tolerant quantum computer require non-Clifford gates such as the $T$-gate – this makes real-time decoding a necessity since syndromes for a patch must be decoded before the application of a non-Clifford gate. We call decodes that must happen before the application of a non-Clifford gate critical decodes, since all syndromes generated up to that point for that logical qubit must be decoded before computation can proceed. Finally, performing non-Clifford operations on logical qubits encoded in quantum Low Distance Parity Check (qLDPC) codes require a critical decode to measure the ancillary system required to transfer the logical qubit to a Surface Code patch (Cross et al., 2024).
@@ 3. Classical Processing Requirements
We now discuss the classical processing requirements for FTQC.
@@ 3.1. Syndrome Generation and Processing Rates
As discussed in Section 2, applying non-Clifford gates requires decoders to be up-to date with the latest syndrome for a logical qubit before computation can proceed. As shown by Terhal (Terhal, 2015, p. 20), if the rate at which syndromes are generated $r_{gen}$ is faster than the rate at which they are processed $r_{proc}$, the time required to execute a workload increases exponentially with the number of non-Clifford $T$ states required by the program (the backlog problem (Terhal, 2015)). Now, if every decoder in the system has an $r_{proc}>r_{gen}$, and every logical qubit has its own decoder, there will be no backlog of syndromes to be decoded and hence no slowdown in the system. Since non-Clifford operations will be very frequent in FTQC systems (Beverland et al., 2022b), ensuring that $r_{proc}$ is consistently greater than $r_{gen}$ is a critical requirement for FTQC systems.
3.2. Why is Real-Time Decoding Necessary?
This requirement for syndromes to be processed faster than they are generated has motivated research to build fast and accurate hardware decoders for the Surface Code (Smith et al., 2023; Ravi et al., 2023; Barber et al., 2023; Vittal et al., 2023; Alavisamani et al., 2024), especially for systems using superconducting qubit architectures due to their fast gate times. While the syndrome processing rates achieved by these decoders are far higher than typical syndrome generation rates achieved today (Acharya et al., 2023; Bluvstein et al., 2023), leaving syndromes undecoded for many successive rounds can be problematic. Figure 4(a) shows how the decoder latency normalized to the number of rounds (latency per round) processed by the decoder can increase exponentially with the number of rounds of undecoded syndromes, especially for $d=7,9$ (circuit-level noise $p=10^{-4}$). The number of rounds was chosen as a multiple of $d$ since it represents the shortest period required for executing a logical operation (Horsman et al., 2012). This slowdown is easily explainable by the fact that errors will accumulate the longer a logical qubit remains undecoded, thus requiring more corrections to be performed. This can result in a significant slowdown, and hence a decrease in $r_{proc}$ leading to higher memory requirements to store undecoded syndromes. However, note that the slowdown in $r_{proc}$ will result in more rounds of error correction required to complete the computation. Figure 4(b) shows how the logical error rate grows slowly with the number of rounds for $p=10^{-3}$, $p=10^{-4}$ respectively. Code distances are selected to achieve a target logical error rate after the $N$ rounds it takes to complete a program (Beverland et al., 2022b; Blunt et al., 2024) – critical decodes delayed exponentially due to a slow $r_{proc}$ will exacerbate the logical error rate, since more rounds would be needed to complete the computation.